# Intelligent-Based Neural Networks and Optimal Control of Fractional Order Ebola Virus Dynamics

### v1

@@ 4.1 Well-Posedness of the Model
Theorem 4.1 (Existence and Uniqueness)
Given any $\alpha\in(0,1]$, and initial conditions
$\mathbf{X}_{0}=(S_{0},V_{0},E_{0},I_{s0},I_{a0},H_{0},D_{0},R_{0})\in\mathbb{R}_{+}^{8}$, the fractional-order system (2) has the unique solution $\mathbf{X}(t)=(S(t),V(t),E(t),I_{s}(t),I_{a}(t),H(t),D(t),R(t))$ for all $t\geq 0$.
Proof.  Define the state vector $\mathbf{X}(t)=(S,V,E,I_{s},I_{a},H,D,R)^{T}$ and express the system as:
@@
The equivalent Volterra integral formulation is:
$\mathbf{X}(t)=\mathbf{X}_{0}+\frac{1}{\Gamma(\alpha)}\int_{0}^{t}(t-s)^{\alpha-1}\mathbf{F}(s,\mathbf{X}(s))ds.$ (3)
Assume the operator $\mathcal{T}:C([0,T],\mathbb{R}^{8})\to C([0,T],\mathbb{R}^{8})$ defined by:
$(\mathcal{T}\mathbf{X})(t)=\mathbf{X}_{0}+\frac{1}{\Gamma(\alpha)}\int_{0}^{t}(t-s)^{\alpha-1}\mathbf{F}(s,\mathbf{X}(s))ds.$
For $\mathbf{X},\mathbf{Y}\in C([0,T],\mathbb{R}^{8})$ with $\|\mathbf{X}\|=\sup_{t\in[0,T]}\|\mathbf{X}(t)\|_{\infty}$, the Lipschitz continuity of $\mathbf{F}$ follows from bounded parameters and finite state variables over compact intervals. Specifically, there exists $L>0$ such that:
$\|\mathbf{F}(t,\mathbf{X})-\mathbf{F}(t,\mathbf{Y})\|_{\infty}\leq L\|\mathbf{X}-\mathbf{Y}\|_{\infty}.$
The contraction estimate yields:
$\|\mathcal{T}\mathbf{X}-\mathcal{T}\mathbf{Y}\|\leq\frac{L\tau^{\alpha}}{\Gamma(\alpha+1)}\|\mathbf{X}-\mathbf{Y}\|.$
Selecting $\tau>0$ such that $\frac{L\tau^{\alpha}}{\Gamma(\alpha+1)}<1$ ensures $\mathcal{T}$ is a contraction on $C([0,\tau],\mathbb{R}^{8})$. Banach’s fixed point theorem guarantees a unique local solution $\mathbf{X}^{*}\in C([0,\tau],\mathbb{R}^{8})$.
Global extension to $0,\infty)$ follows from the linear growth condition $\|\mathbf{F}(t,\mathbf{X})\|_{\infty}\leq M(1+\|\mathbf{X}\|_{\infty})$ for some $M>0$, which precludes finite-time blow-up. Iterative continuation establishes global existence and uniqueness. $\square$
@@ 4.2 Invariant Region and Attractivity
The fractional-order Ebola system (2) evolves within the biologically feasible region:
@@
Theorem 4.2 (Invariance and Attractivity)
The region $\Omega$ is:
@@
    Positively invariant under (2),
@@
Proof.  (i) Positivity and Invariance: Applying the fractional positivity lemma [29], we verify boundary behavior:
${}^{C}D_{t}^{\alpha}S|_{S=0}=\Lambda+\omega V\geq 0,\quad{}^{C}D_{t}^{\alpha}V|_{V=0}=vS\geq 0,$
@@
Non-negativity of all boundary fluxes ensures $\mathbb{R}_{+}^{8}$ is invariant. (ii) Boundedness and Attractivity: For the living population $N_{L}$:
$^{C}D_{t}^{\alpha}N_{L}=\Lambda-\mu N_{L}-(\delta_{s}I_{s}+\delta_{h}H)\leq\Lambda-\mu N_{L}.$
Laplace transformation yields:
$N_{L}(s)\leq\frac{\Lambda}{s(s^{\alpha}+\mu)}+N_{L}(0)\frac{s^{\alpha-1}}{s^{\alpha}+\mu},$
with inverse transform:
@@
As $t\to\infty$, $E_{\alpha,1}(-\mu t^{\alpha})\to 0$ and $t^{\alpha}E_{\alpha,\alpha+1}(-\mu t^{\alpha})\to 1/\mu$, giving:
@@
For deceased individuals:
@@
4.3 EVD-Free Equilibrium
The state wherein a population exists without any infections of the Ebola virus is referred to the Ebola-free equilibrium (DFE), at which point all infected compartments have completely dissipated:
@@ 4.3 EVD-Free Equilibrium
The system (2) reduces to:
$0$ $=\Lambda-q_{0}S+\omega V,$ (5)
$0$ $=vS-q_{1}V,$ (6)
$0$ $=-\mu R.$ (7)
From (7), $R^{*}=0$. Equations (5) and (6) form the linear system:
$q_{0}S-\omega V$ $=\Lambda,$ (8)
$-vS+q_{1}V$ $=0.$ (9)
@@
ensuring a unique solution. Applying Cramer’s rule:
@@
Thus, the disease-free equilibrium is:
$\mathcal{E}_{0}=\left(\frac{\Lambda q_{1}}{\mu(q_{1}+v)},\frac{\Lambda v}{\mu(q_{1}+v)},0,0,0,0,0,0\right).$
This biologically consistent DFE provides the foundation for stability analysis and reproduction number derivation.
[…]
$\mathcal{R}_{0}$ is computed via the next generation matrix method as expressed in [55]. In the case of the infected subsystem $X=(E,I_{s},I_{a},H,D)^{T}$, the new infection ($F$) and transition ($V$) matrices are as follows:
$F=\begin{pmatrix}0&\beta\Psi&\beta\eta_{a}\Psi&0&\beta\eta_{d}\Psi\\ 0&0&0&0&0\\ 0&0&0&0&0\\ 0&0&0&0&0\\ 0&0&0&0&0\end{pmatrix},\quad V=\begin{pmatrix}q_{2}&0&0&0&0\\ -p\sigma&q_{3}&0&0&0\\ -(1-p)\sigma&0&q_{4}&0&0\\ 0&-h_{s}&0&q_{5}&0\\ 0&-\delta_{s}&0&-\delta_{h}&q_{6}\end{pmatrix}$
where $\Psi=\frac{S^{*}+(1-\varepsilon)V^{*}}{N^{*}}$ with $N^{*}=\frac{\Lambda}{\mu}$.
The inverse transition matrix is:
$V^{-1}=\begin{pmatrix}\frac{1}{q_{2}}&0&0&0&0\\ \frac{p\sigma}{q_{2}q_{3}}&\frac{1}{q_{3}}&0&0&0\\ \frac{(1-p)\sigma}{q_{2}q_{4}}&0&\frac{1}{q_{4}}&0&0\\ \frac{p\sigma h_{s}}{q_{2}q_{3}q_{5}}&\frac{h_{s}}{q_{3}q_{5}}&0&\frac{1}{q_{5}}&0\\ \frac{p\sigma(\delta_{s}q_{5}+\delta_{h}h_{s})}{q_{2}q_{3}q_{5}q_{6}}&\frac{\delta_{s}q_{5}+\delta_{h}h_{s}}{q_{3}q_{5}q_{6}}&0&\frac{\delta_{h}}{q_{5}q_{6}}&\frac{1}{q_{6}}\end{pmatrix}$
The spectral radius of the next generation matrix is $K=FV^{-1}$ yields:
$\mathcal{R}_{0}=\beta\Psi\left[\frac{p\sigma}{q_{2}}\left(\frac{1}{q_{3}}+\frac{\eta_{d}(\delta_{s}q_{5}+\delta_{h}h_{s})}{q_{3}q_{5}q_{6}}\right)+\frac{\eta_{a}(1-p)\sigma}{q_{2}q_{4}}\right]$
Substituting $\Psi$ provides the explicit form:
$\mathcal{R}_{0}=\frac{\beta\sigma[S^{*}+(1-\varepsilon)V^{*}]}{(\Lambda/\mu)q_{2}}\left[\frac{p}{q_{3}}\left(1+\frac{\eta_{d}(\delta_{s}q_{5}+\delta_{h}h_{s})}{q_{5}q_{6}}\right)+\frac{\eta_{a}(1-p)}{q_{4}}\right]$
This threshold quantity governs disease invasion and extinction dynamics.
@@
4.5 Critical Vaccination Coverage
The critical vaccination coverage defines the minimum population immunity required for disease elimination, occurring when $\mathcal{R}_{0}=1$. Beginning with the reproduction number:
$\mathcal{R}_{0}(v)=\mathcal{R}_{0}(0)\cdot\frac{q_{1}+(1-\varepsilon)v}{q_{1}+v}$
where $\mathcal{R}_{0}(0)=K$ represents the vaccination-free reproduction number. Setting $\mathcal{R}_{0}(v_{c})=1$ yields:
$1=\mathcal{R}_{0}(0)\cdot\frac{q_{1}+(1-\varepsilon)v_{c}}{q_{1}+v_{c}}$
Solving for the critical vaccination rate $v_{c}$:
$q_{1}+v_{c}$ $=\mathcal{R}_{0}(0)[q_{1}+(1-\varepsilon)v_{c}]$
$v_{c}-\mathcal{R}_{0}(0)(1-\varepsilon)v_{c}$ $=\mathcal{R}_{0}(0)q_{1}-q_{1}$
$v_{c}[1-\mathcal{R}_{0}(0)(1-\varepsilon)]$ $=q_{1}[\mathcal{R}_{0}(0)-1]$
$v_{c}=\frac{q_{1}[\mathcal{R}_{0}(0)-1]}{1-(1-\varepsilon)\mathcal{R}_{0}(0)}$
The critical vaccination coverage $p_{c}$, representing the proportion vaccinated at equilibrium, is:
$p_{c}=\frac{v_{c}}{q_{1}+v_{c}}=\frac{\mathcal{R}_{0}(0)-1}{\varepsilon\mathcal{R}_{0}(0)}$
Thus, the herd immunity threshold for Ebola elimination is:
$p_{c}=\max\left(0,1-\frac{1}{\varepsilon\mathcal{R}_{0}(0)}\right)$
This establishes the minimal vaccination coverage required to interrupt disease transmission.
Figure 4: 3D surface: Vaccination and waning immunity impact on $\mathcal{R}_{0}$. The surface shows how the basic reproduction number varies with vaccination rate ($v$) and waning immunity rate ($\omega$).
4.6 Local Stability Analysis of the EVD-Free Equilibrium
@@
When $\mathcal{R}_{0}<1$, the EVD-free equilibrium $\mathcal{E}_{0}$ of the system (2) is achieved and remains locally stable; however, it becomes unsteady and $\mathcal{R}_{0}>1$ will prove it to be unstable.
Proof.  The Jacobian matrix evaluated at $\mathcal{E}_{0}$ takes the block triangular form:
$J(\mathcal{E}_{0})=\begin{pmatrix}J_{11}&J_{12}\\ 0&J_{22}\end{pmatrix}$
where $J_{11}$ corresponds to the uninfected compartments $(S,V,R)$ and $J_{22}$ to the infected compartments $(E,I_{s},I_{a},H,D)$. The uninfected subsystem matrix is:
$J_{11}=\begin{pmatrix}-(v+\mu)&\omega&0\\ v&-(\mu+\omega)&0\\ 0&0&-\mu\end{pmatrix}$
with eigenvalues $\lambda_{1}=-\mu$, $\lambda_{2}=-\mu$, $\lambda_{3}=-(\mu+v+\omega)$. All eigenvalues satisfy $|\arg(\lambda_{i})|=\pi>\frac{\alpha\pi}{2}$ for $\alpha\in(0,1]$, confirming stability of the uninfected subsystem. The infected subsystem Jacobian is:
$J_{22}=\begin{pmatrix}-q_{2}&\beta\Psi&\beta\eta_{a}\Psi&0&\beta\eta_{d}\Psi\\ p\sigma&-q_{3}&0&0&0\\ (1-p)\sigma&0&-q_{4}&0&0\\ 0&h_{s}&0&-q_{5}&0\\ 0&\delta_{s}&0&\delta_{h}&-q_{6}\end{pmatrix}$
where $\Psi=\frac{S^{*}+(1-\varepsilon)V^{*}}{N^{*}}$. We decompose $J_{22}=F-V$, where $F$ and $V$ are the next-generation matrices previously defined. The stability of $J_{22}$ is determined by the spectral radius $\rho(FV^{-1})=\mathcal{R}_{0}$. According to the fractional Routh-Hurwitz criteria and properties of Metzler matrices, when $\mathcal{R}_{0}<1$, all eigenvalues of $J_{22}$ have negative real parts and satisfy the fractional stability condition $|\arg(\lambda)|>\frac{\alpha\pi}{2}$.
Conversely, when $\mathcal{R}_{0}>1$, the spectral abscissa of $J_{22}$ is positive, indicating at least one eigenvalue with positive real part. By the fractional stability theorem, this establishes instability of the disease-free equilibrium. $\square$
4.7 Global Stability of EVD-Free Equilibrium
@@
When $\mathcal{R}_{0}<1$, the EVD-free equilibrium $\mathcal{E}_{0}$ of the system (2) is globally asymptotically stable in $\Omega$.
Proof.  Consider the function for the infected compartments already introduced in the above section.
$L(E,I_{s},I_{a},H,D)=E+AI_{s}+BI_{a}+CH+FD,$
where the parameters A, B, C, and F are positive and not arbitrary. To investigate the function, we follow the trajectories described in the system and compute the Caputo derivative along each trajectory:
${}^{C}D_{t}^{\alpha}L=$ $\left[[\frac{c\beta[S+(1-\varepsilon)V](I_{s}+\eta_{a}I_{a}+\eta_{d}D)}{N}]-q_{2}E\right]$
$+A\left(p\sigma E-q_{3}I_{s}\right)+B((1-p)\sigma E-q_{4}I_{a})$
$+C\left(h_{s}I_{s}-q_{5}H\right)+F(\delta_{s}I_{s}+\delta_{h}H-q_{6}D)$
Rearranging terms by compartment:
@@
$+\left[\frac{c\beta[S+(1-\varepsilon)V]}{N}-Aq_{3}+Ch_{s}+F\delta_{s}\right]I_{s}$
$+\left[\frac{c\beta\eta_{a}[S+(1-\varepsilon)V]}{N}-Bq_{4}\right]I_{a}$
$+\left[-Cq_{5}+F\delta_{h}\right]H$
$+\left[\frac{c\beta\eta_{d}[S+(1-\varepsilon)V]}{N}-Fq_{6}\right]D$
We select coefficients to eliminate positive terms:
$C$ $=\frac{F\delta_{h}}{q_{5}},\quad B=\frac{\beta\eta_{a}\Psi}{q_{4}},\quad F=\frac{\beta\eta_{d}\Psi}{q_{6}}$
$A$ $=\frac{\beta\Psi}{q_{3}}+\frac{Ch_{s}}{q_{3}}+\frac{F\delta_{s}}{q_{3}}=\frac{\beta\Psi}{q_{3}}+\frac{F\delta_{h}h_{s}}{q_{3}q_{5}}+\frac{F\delta_{s}}{q_{3}}$
Using the bound $\frac{S+(1-\varepsilon)V}{N}\leq\Psi$ valid throughout $\Omega$, we obtain:
@@
$+\beta\Psi\left[1-1\right]I_{s}+\beta\eta_{a}\Psi\left[1-1\right]I_{a}+\beta\eta_{d}\Psi\left[1-1\right]D$
Substituting $F=\frac{\beta\eta_{d}\Psi}{q_{6}}$ yields:
$^{C}D_{t}^{\alpha}L\leq q_{2}\left[-1+\mathcal{R}_{0}\right]E$
For $\mathcal{R}_{0}\leq 1$, we have $^{C}D_{t}^{\alpha}L\leq 0$ throughout $\Omega$, with equality if and only if $E=0$. When $E=0$, the system dynamics ensure:
${}^{C}D_{t}^{\alpha}I_{s}$ $=-q_{3}I_{s}\Rightarrow I_{s}\to 0$
${}^{C}D_{t}^{\alpha}I_{a}$ $=-q_{4}I_{a}\Rightarrow I_{a}\to 0$
${}^{C}D_{t}^{\alpha}H$ $=h_{s}I_{s}-q_{5}H\Rightarrow H\to 0$
${}^{C}D_{t}^{\alpha}D$ $=\delta_{s}I_{s}+\delta_{h}H-q_{6}D\Rightarrow D\to 0$
Thus, the maximal invariant set where $^{C}D_{t}^{\alpha}L=0$ is precisely $\mathcal{E}_{0}$. Application of the fractional LaSalle invariance principle establishes global asymptotic stability when $\mathcal{R}_{0}\leq 1$. For $\mathcal{R}_{0}>1$, instability follows from local analysis, while uniform persistence can be demonstrated via established methods in fractional dynamical systems [48]. $\square$

### v3

@@ 4.1 Well-Posedness of the Model
Theorem 4.1
Given any $\alpha\in(0,1]$, and initial conditions ae $\mathbf{X}_{0}=(S_{0},V_{0},E_{0},I_{s0},I_{a0},H_{0},D_{0},R_{0})\in\mathbb{R}_{+}^{8}$, the fractional-order model (2) has the unique solution $\mathbf{X}(t)=(S(t),V(t),E(t),I_{s}(t),I_{a}(t),H(t),D(t),R(t))$ for all $t\geq 0$.
Proof.  Define the state vector $\mathbf{X}(t)=(S,V,E,I_{s},I_{a},H,D,R)^{T}$ and express the system as
@@
The equivalent Volterra integral formulation is
$\mathbf{X}(t)=\mathbf{X}_{0}+\frac{1}{\Gamma(\alpha)}\int_{0}^{t}(t-s)^{\alpha-1}\mathbf{F}(s,\mathbf{X}(s))ds,$ (3)
Assume the operator $\mathcal{T}:C([0,T],\mathbb{R}^{8})\to C([0,T],\mathbb{R}^{8})$ defined by
$(\mathcal{T}\mathbf{X})(t)=\mathbf{X}_{0}+\frac{1}{\Gamma(\alpha)}\int_{0}^{t}(t-s)^{\alpha-1}\mathbf{F}(s,\mathbf{X}(s))ds,$
For $\mathbf{X},\mathbf{Y}\in C([0,T],\mathbb{R}^{8})$ with $\|\mathbf{X}\|=\sup_{t\in[0,T]}\|\mathbf{X}(t)\|_{\infty}$, the Lipschitz continuity of $\mathbf{F}$ follows from bounded parameters and finite state variables over compact intervals. Specifically, there exists $L>0$ such that
$\|\mathbf{F}(t,\mathbf{X})-\mathbf{F}(t,\mathbf{Y})\|_{\infty}\leq L\|\mathbf{X}-\mathbf{Y}\|_{\infty},$ (4)
The contraction estimate provides that
$\|\mathcal{T}\mathbf{X}-\mathcal{T}\mathbf{Y}\|\leq\frac{L\tau^{\alpha}}{\Gamma(\alpha+1)}\|\mathbf{X}-\mathbf{Y}\|.$ (5)
[…]
@@ 4.2 Invariant Region and Attractivity
The fractional-order Ebola model (2) evolves within the biologically feasible region is
@@
Theorem 4.2
The region $\Omega$ is defined as
@@
    Positively invariant under model (2),
@@
Proof.  Applying the fractional positivity lemma from [21], we verify boundary behavior as,
${}^{C}D_{t}^{\alpha}S|_{S=0}=\Lambda+\omega V\geq 0,\quad{}^{C}D_{t}^{\alpha}V|_{V=0}=vS\geq 0,$ (7)
@@
Non-negativity of all boundary fluxes ensures $\mathbb{R}_{+}^{8}$ is invariant.
The total living population, defined as $N_{L}(t)=S(t)+V(t)+E(t)+I_{s}(t)+I_{a}(t)+H(t)+R(t)$, evolves according to
$^{C}D_{t}^{\alpha}N_{L}=\Lambda-\mu N_{L}-(\delta_{s}I_{s}+\delta_{h}H)\leq\Lambda-\mu N_{L},$ (8)
Applying the Laplace transform to the inequality $^{C}D_{t}^{\alpha}N_{L}\leq\Lambda-\mu N_{L}$ gives
$N_{L}(s)\leq\frac{\Lambda}{s(s^{\alpha}+\mu)}+N_{L}(0)\frac{s^{\alpha-1}}{s^{\alpha}+\mu},$ (9)
Applying the inverse Laplace transform yields
@@
Taking the limit as $t\to\infty$, we use the asymptotic behaviors $E_{\alpha,1}(-\mu t^{\alpha})\to 0$ and $t^{\alpha}E_{\alpha,\alpha+1}(-\mu t^{\alpha})\to 1/\mu$,
which gives the bound
@@
For deceased individuals, we derive from the differential inequality
@@
4.3 Ebola-Free Equilibrium
The state in which the population is completely free of Ebola virus infection is known as the disease-free equilibrium (DFE). At this equilibrium, all infected compartments vanish, meaning
@@ 4.3 Ebola-Free Equilibrium
The system (2) then reduces to
$0$ $=\Lambda-q_{0}S+\omega V,$
$0$ $=vS-q_{1}V,$
$0$ $=-\mu R.$
From the third equation, we obtain $R^{*}=0$. The first two equations form the linear system
$q_{0}S-\omega V$ $=\Lambda,$
$-vS+q_{1}V$ $=0.$
@@
which guarantees a unique solution. Applying Cramer’s rule yields
@@
Thus, the disease-free equilibrium is given by
$\mathcal{E}_{0}=\left(\frac{\Lambda q_{1}}{\mu(q_{1}+v)},\frac{\Lambda v}{\mu(q_{1}+v)},0,0,0,0,0,0\right).$ (10)
This biologically consistent DFE serves as the basis for stability analysis and the derivation of the basic reproduction number.
4.4 Basic Reproduction Number $\mathcal{R}_{0}$
The basic reproduction number $\mathcal{R}_{0}$ is computed using the next-generation matrix method [53]. Considering the infected subsystem $X=(E,I_{s},I_{a},H,D)^{T}$, we define the new infection matrix $F$ and the transition matrix $V$ as
$F=\begin{pmatrix}0&\beta\Psi&\beta\eta_{a}\Psi&0&\beta\eta_{d}\Psi\\ 0&0&0&0&0\\ 0&0&0&0&0\\ 0&0&0&0&0\\ 0&0&0&0&0\end{pmatrix},\qquad V=\begin{pmatrix}q_{2}&0&0&0&0\\ -p\sigma&q_{3}&0&0&0\\ -(1-p)\sigma&0&q_{4}&0&0\\ 0&-h_{s}&0&q_{5}&0\\ 0&-\delta_{s}&0&-\delta_{h}&q_{6}\end{pmatrix}$,
where $\Psi=\frac{S^{*}+(1-\varepsilon)V^{*}}{N^{*}}$ and $N^{*}=\frac{\Lambda}{\mu}$. The inverse of the transition matrix is
$V^{-1}=\begin{pmatrix}\frac{1}{q_{2}}&0&0&0&0\\ \frac{p\sigma}{q_{2}q_{3}}&\frac{1}{q_{3}}&0&0&0\\ \frac{(1-p)\sigma}{q_{2}q_{4}}&0&\frac{1}{q_{4}}&0&0\\ \frac{p\sigma h_{s}}{q_{2}q_{3}q_{5}}&\frac{h_{s}}{q_{3}q_{5}}&0&\frac{1}{q_{5}}&0\\ \frac{p\sigma(\delta_{s}q_{5}+\delta_{h}h_{s})}{q_{2}q_{3}q_{5}q_{6}}&\frac{\delta_{s}q_{5}+\delta_{h}h_{s}}{q_{3}q_{5}q_{6}}&0&\frac{\delta_{h}}{q_{5}q_{6}}&\frac{1}{q_{6}}\end{pmatrix}$.
The next-generation matrix is $K=FV^{-1}$, and its spectral radius gives the basic reproduction number
$\mathcal{R}_{0}=\beta\Psi\left[\frac{p\sigma}{q_{2}}\left(\frac{1}{q_{3}}+\frac{\eta_{d}(\delta_{s}q_{5}+\delta_{h}h_{s})}{q_{3}q_{5}q_{6}}\right)+\frac{\eta_{a}(1-p)\sigma}{q_{2}q_{4}}\right].$ (11)
Substituting the expression for $\Psi$ yields the explicit form
$\mathcal{R}_{0}=\frac{\beta\sigma[S^{*}+(1-\varepsilon)V^{*}]}{(\Lambda/\mu)q_{2}}\left[\frac{p}{q_{3}}\left(1+\frac{\eta_{d}(\delta_{s}q_{5}+\delta_{h}h_{s})}{q_{5}q_{6}}\right)+\frac{\eta_{a}(1-p)}{q_{4}}\right].$ (12)
This threshold quantity governs the disease invasion and extinction dynamics.
@@ 4.4 Basic Reproduction Number $\mathcal{R}_{0}$
4.5 Local Stability Analysis of the Disease-Free Equilibrium
@@
The disease-free equilibrium $\mathcal{E}_{0}$ of the model (2) is locally asymptotically stable when $\mathcal{R}_{0}<1$, and unstable when $\mathcal{R}_{0}>1$.
Proof.  The Jacobian matrix evaluated at $\mathcal{E}_{0}$ exhibits a block triangular structure
$J(\mathcal{E}_{0})=\begin{pmatrix}J_{11}&J_{12}\\ 0&J_{22}\end{pmatrix}$,
where $J_{11}$ corresponds to the uninfected compartments $(S,V,R)$ and $J_{22}$ corresponds to the infected compartments $(E,I_{s},I_{a},H,D)$.
For the uninfected subsystem, the matrix is
$J_{11}=\begin{pmatrix}-(v+\mu)&\omega&0\\ v&-(\mu+\omega)&0\\ 0&0&-\mu\end{pmatrix}$.
The eigenvalues of $J_{11}$ are $\lambda_{1}=-\mu$, $\lambda_{2}=-\mu$, and $\lambda_{3}=-(\mu+v+\omega)$. For $\alpha\in(0,1]$, all eigenvalues satisfy $|\arg(\lambda_{i})|=\pi>\alpha\pi/2$, ensuring the stability of the uninfected subsystem.
For the infected subsystem, the Jacobian matrix takes the form
$J_{22}=\begin{pmatrix}-q_{2}&\beta\Psi&\beta\eta_{a}\Psi&0&\beta\eta_{d}\Psi\\ p\sigma&-q_{3}&0&0&0\\ (1-p)\sigma&0&-q_{4}&0&0\\ 0&h_{s}&0&-q_{5}&0\\ 0&\delta_{s}&0&\delta_{h}&-q_{6}\end{pmatrix}$,
where $\Psi=\frac{S^{*}+(1-\varepsilon)V^{*}}{N^{*}}$.
We observe that $J_{22}$ can be decomposed as $J_{22}=F-V$, where $F$ and $V$ are the next-generation matrices defined previously. The stability of $J_{22}$ is governed by the spectral radius $\rho(FV^{-1})=\mathcal{R}_{0}$. According to the fractional Routh-Hurwitz criteria and the properties of Metzler matrices, when $\mathcal{R}_{0}<1$, all eigenvalues of $J_{22}$ have negative real parts and satisfy the fractional stability condition $|\arg(\lambda)|>\alpha\pi/2$.
Conversely, when $\mathcal{R}_{0}>1$, the spectral abscissa of $J_{22}$ is positive, indicating at least one eigenvalue with a positive real part. By the fractional stability theorem, this establishes the instability of the disease-free equilibrium. $\square$
4.6 Global Stability of the Disease-Free Equilibrium
@@
If $\mathcal{R}_{0}<1$, the disease-free equilibrium $\mathcal{E}_{0}$ of the model (2) is globally asymptotically stable in the invariant region $\Omega$.
Proof.  We construct a Lyapunov function for the infected compartments based on the approach outlined in the previous section
$L(E,I_{s},I_{a},H,D)=E+AI_{s}+BI_{a}+CH+FD,$ (13)
where $A$, $B$, $C$, and $F$ are positive constants to be determined.
Following the system trajectories, we compute the Caputo derivative along solutions
${}^{C}D_{t}^{\alpha}L$ $=\left[\frac{c\beta[S+(1-\varepsilon)V](I_{s}+\eta_{a}I_{a}+\eta_{d}D)}{N}-q_{2}E\right]$
$\quad+A\left(p\sigma E-q_{3}I_{s}\right)+B\left((1-p)\sigma E-q_{4}I_{a}\right)$
$\quad+C\left(h_{s}I_{s}-q_{5}H\right)+F\left(\delta_{s}I_{s}+\delta_{h}H-q_{6}D\right).$
Rearranging terms by compartment yields
@@
$\quad+\left[\frac{c\beta[S+(1-\varepsilon)V]}{N}-Aq_{3}+Ch_{s}+F\delta_{s}\right]I_{s}$
$\quad+\left[\frac{c\beta\eta_{a}[S+(1-\varepsilon)V]}{N}-Bq_{4}\right]I_{a}$
$\quad+\left[-Cq_{5}+F\delta_{h}\right]H$
$\quad+\left[\frac{c\beta\eta_{d}[S+(1-\varepsilon)V]}{N}-Fq_{6}\right]D.$
We now select the coefficients to eliminate positive terms from the $I_{s}$, $I_{a}$, $H$, and $D$ components
$C$ $=\frac{F\delta_{h}}{q_{5}},\quad B=\frac{\beta\eta_{a}\Psi}{q_{4}},\quad F=\frac{\beta\eta_{d}\Psi}{q_{6}},$
$A$ $=\frac{\beta\Psi}{q_{3}}+\frac{Ch_{s}}{q_{3}}+\frac{F\delta_{s}}{q_{3}}=\frac{\beta\Psi}{q_{3}}+\frac{F\delta_{h}h_{s}}{q_{3}q_{5}}+\frac{F\delta_{s}}{q_{3}},$ (14)
where $\Psi=\frac{S^{*}+(1-\varepsilon)V^{*}}{N^{*}}$.
Using the inequality $\frac{S+(1-\varepsilon)V}{N}\leq\Psi$, which holds throughout $\Omega$, we obtain the estimate
@@
$\quad+\beta\Psi\left[1-1\right]I_{s}+\beta\eta_{a}\Psi\left[1-1\right]I_{a}+\beta\eta_{d}\Psi\left[1-1\right]D.$
Substituting $F=\frac{\beta\eta_{d}\Psi}{q_{6}}$ and simplifying leads to the key inequality
$^{C}D_{t}^{\alpha}L\leq q_{2}\left(-1+\mathcal{R}_{0}\right)E.$ (15)
When $\mathcal{R}_{0}\leq 1$, we have $^{C}D_{t}^{\alpha}L\leq 0$ throughout $\Omega$, with equality holding only when $E=0$. If $E=0$, the system dynamics imply
${}^{C}D_{t}^{\alpha}I_{s}$ $=-q_{3}I_{s}\quad\Rightarrow\quad I_{s}\to 0,$
${}^{C}D_{t}^{\alpha}I_{a}$ $=-q_{4}I_{a}\quad\Rightarrow\quad I_{a}\to 0,$
${}^{C}D_{t}^{\alpha}H$ $=h_{s}I_{s}-q_{5}H\quad\Rightarrow\quad H\to 0,$
${}^{C}D_{t}^{\alpha}D$ $=\delta_{s}I_{s}+\delta_{h}H-q_{6}D\quad\Rightarrow\quad D\to 0.$
Therefore, the maximal invariant set where $^{C}D_{t}^{\alpha}L=0$ is exactly the disease-free equilibrium $\mathcal{E}_{0}$. By the fractional LaSalle invariance principle, $\mathcal{E}_{0}$ is globally asymptotically stable in $\Omega$ when $\mathcal{R}_{0}\leq 1$.
For the case $\mathcal{R}_{0}>1$, instability follows from the local analysis, while uniform persistence can be established using standard methods for fractional dynamical systems. $\square$